1. A rugby player runs 100. m from his own goal line to the opposing team’s goal line, returning to the halfway line, all in 18.0 s. Calculate (a) his average speed, and (b) the magnitude of his average velocity.
1
Expert's answer
2018-08-08T15:44:08-0400
The velocity
v(t)=x^' (t)=2+6t+12t^2
v(2)=2+6×2+12×2^2=62 m/s
The acceleration
a(t)=v^' (t)=6+24t
a(3)=6+24×3=78 m/s^2